### v1

# SOP-Bench: Complex Industrial SOPs for Evaluating LLM Agents

Abstract
Large Language Models (LLMs) have demonstrated impressive general-purpose reasoning, language understanding and problem-solving abilities. However, they continue to fall short in executing complex, long-horizon workflows that require strict adherence to Standard Operating Procedures (SOPs), a critical requirement for real-world industrial automation. Despite this, there is a lack of publicly available benchmarks that capture the complexity, structure and domain-specific nuances of industrial SOPs. To address this gap, we present three main contributions. First, we introduce a novel synthetic data generation framework for creating realistic, industry-grade SOPs that rigorously test the planning, reasoning, and tool-use capabilities of LLM-based agents. Second, using this framework, we develop SOP-Bench, a benchmark comprising over 1,800 tasks across 10 industrial domains, each coupled with APIs, tool interfaces, and human-validated test cases. Third, we evaluate two prominent agent architectures: Function-Calling Agent and ReAct Agent, on SOP-Bench, and observe average task success rates of only 27% and 48%, respectively. Remarkably, in tasks where the tools registry is much larger than what is needed for task completion, the agents invoke an incorrect tool nearly 100% of the time. Our findings underscore a substantial gap between current LLM’s agentic capabilities and the demands of real-world SOPs’ successful automations.
Furthermore, performance varies considerably by task and domain, which highlights the need for domain-specific benchmarking and careful architectural selection prior to agent deployment. SOP-Bench is publicly available at http://sop-bench.s3-website-us-west-2.amazonaws.com/. We are also releasing all the prompts that underpin the complete data generation framework to support creation of new synthetic SOPs for domain-specific evaluation. We invite the community to extend SOP-Bench by contributing the SOPs from their own industrial domains.
1 Introduction
Standard Operating Procedures (SOPs) are essential to consistent and reliable operations across industries, encoding domain expertise and compliance rules into structured workflows [4, 5]. With recent advances in large language models (LLMs), there is growing interest in automating SOP execution using LLM-based agents [25, 30, 2]. However, deploying such agents in production poses challenges that are not captured by existing benchmarks like Gorilla [16], API-Bank [9], ComplexBench [27].
Criteria SLUHN-SOP UNTHS-SOP
Length Long (7 pgs) Short (3 pgs)
Instruction ambiguity Low Moderate
Complexity of steps High Low to moderate
Implicit knowledge needed High Low to moderate
Modularity High Moderate
Table 1: Highlighting the diveristy of SOPs in industry, assessed by OpenAI’s GPT-4-turbo [15] and validated by human associates with extensive cross-domain SOP review experience.
These challenges are particularly acute in knowledge operations like verification services, support ticket triaging, and customer support, where SOPs involve ambiguous language, implicit domain knowledge, and complex decision trees [21]. For example, consider the two patient registration SOPs in appendix E, one from St. Luke’s University Health Network (SLUHN) [^1] https://www.slhn.org/-/media/slhn/Research/File/PDF/SOPs-and-Policies/SOP-302Patient-Registration-and-Ongoing-Subject-MgmtV30CLEAN71216.ashx, and one from University of North Texas Health Science Center (UNTHS)  [^2] https://www.unthsc.edu/wp-content/uploads/sites/23/Patient_Registration.pdf. Table 1 shows how different two SOPs of the same domain can be.
Generally speaking, industrial knowledge operations SOPs, spanning domains like customer service, content moderation, and data annotation, present challenges distinct from traditional process automation tasks [20]. Unlike manufacturing processes, these operations require complex information synthesis and context-dependent decisions in ambiguous situations, with content moderators and annotators navigating evolving guidelines and subjective classifications [21]. These operations frequently encounter edge cases that challenge policy interpretations [6], making them particularly difficult for existing LLM-based automation approaches that assume well-defined success criteria and stable procedures. See figure 1 for an example of an instruction that can be easier for humans to follow but could be ambiguous for LLMs.
While recent advances in LLMs have enabled sophisticated instruction following and tool use [26, 30], automating SOP execution poses unique challenges that existing benchmarks fail to capture. Current evaluation frameworks primarily focus on isolated capabilities—tool use [19], planning [30], or instruction following [26], using clean, synthetic prompts that sidestep the messiness of real-world procedures.
Figure 1: Example of instructions that can be simple for humans but confusing for LLMs: In this excerpt of page 2 of the UNHTS-SOP (appendix E) we can see that steps 4 and 6 both have instructions on verifying the insurance, but does not have explicit information on how to verify. Moreover, it is not clear why there are two steps of insurance verifications. For human associates with implicit knowledge of the domains this might not be as ambiguous as for LLMs. In our experience, this is a common occurrence in industrial SOPs.
To address the gap, we introduce SOP-Bench, a synthetic yet realistic benchmark designed to evaluate LLM agents on SOP-style tasks. SOP-Bench spans 1,800+ tasks across ten domains - including seller appeals triage, and hazardous goods classification, and more. See appendix B for details. The SOPs are paired with synthetic APIs, executable tools, and human-validated test cases. SOP-Bench is designed to assess how agents handle complex control flows, edge cases, and interdependent steps characteristic of industrial SOPs.
We make three contributions: (1) A framework for generating realistic, domain-rich SOP datasets; (2) A benchmark that mirrors industrial SOP complexity while supporting agent evaluation; and (3) Empirical evidence showing that some of the most popular agent paradigms, Function-Calling [13] and ReAct [30], achieve average task success rates of only 27% and 48% respectively, with wide variation across task types, while achieving higher than 90% on tasks in ToolBench [28].
Overall, our goal is to establish a standardized benchmark for evaluating AI agents’ readiness for SOP execution. We encourage the community to use SOP-Bench to exhaustively test different agent paradigms based on different pre-trained models. We are also releasing all prompts and the full data generation framework to enable users to create new synthetic SOPs for domain-specific evaluation and contribute back to SOP-Bench, so that the whole community can benefit from an increasingly larger and more diverse evaluation set.
2 Related Work
LLM Agents for Planning and Tool Use. LLMs have enabled the emergence of AI agents capable of task planning, reasoning, and tool use. Architectures such as ReAct [30], AutoGPT [29], and LangChain [2] have demonstrated autonomous decision-making via multi-step reasoning and API calls. ToolChain [31] and ToolLLM [18] extend this by integrating dynamic tool execution. However, these systems are mostly evaluated in highly simplistic settings, limiting their applicability to real-world operational workflows.
Benchmark
Inst.
Follow
Multi
Step
Tool
Use
Real
Ambig.
Error
Hand.
Work
Flow
Ind.
Ctx.
AlpacaEval [7] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
FollowEval [8] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
ComplexBench [27] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
InFoBench [17] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
Gorilla [16] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
API-Bank [9] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
ToolBench [19] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
AgentBench [12] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
PlanBench [23] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
ALFWorld [22] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
SOP-Bench [Ours] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 2: Comparison of capabilities across different benchmarks. (Column abbreviations: Inst. Follow = Instruction Following (basic ability to follow instructions); Multi Step = Multi-step Tasks (handling sequential, dependent tasks); Tool Use = Tool/API Integration (ability to use external tools and APIs); Real Ambig. = Real-world Ambiguity (handling unclear or incomplete instructions); Error Hand. = Error Handling (ability to handle and recover from errors); Work Flow = Workflow Dependencies (managing inter-task dependencies); Ind. Ctx. = Industrial Context (relevance to real-world industrial settings)).
Instruction Following and Benchmarking. Instruction-following ability has been benchmarked through datasets like SUPER-NATURAL-INSTRUCTIONS [26], AlpacaEval [10], and FollowEval [8]. More recent multi-step benchmarks like ComplexBench [27] and InFoBench [17] explore instruction complexity, but typically use machine-formatted inputs that omit the ambiguity and variability of human-authored SOPs.
Tool Use and API-Centric Benchmarks. Benchmarks like Gorilla [16], API-Bank [9], BENCHAGENTS  [1] and ToolBench [19] assess tool invocation and API usage capabilities. However, these focus on isolated tool interactions with minimal procedural context. In contrast, SOP execution involves coordinated tool use across dependent steps, often requiring error handling and state tracking.
Agent Evaluation Frameworks. AgentBench [12] and PlanBench [23] offer general frameworks to evaluate LLM agents on planning and tool use. Yet, their task settings are narrow in scope and lack the procedural structure, conditional logic, and real-world ambiguity typical of industrial SOPs. Similarly, embodied agents are evaluated in ALFWorld [22] and BabyAI [3], but those environments differ fundamentally from workflow automation tasks.
SOP Automation and Business Process Modeling. Traditional business process automation relies on rule-based systems and formal process modeling languages [5, 11, 24]. While effective, these require manual formalization of procedures, making them brittle and labor-intensive. Prior work on SOP formalization in organizational contexts [4] has not yet translated into LLM-native solutions.
How SOP-Bench Differs SOP-Bench addresses these gaps by offering (1) realistic SOP-style instructions, (2) diverse domain coverage, (3) structured APIs for execution, and (4) evaluation protocols grounded in real-world complexity. Compared to benchmarks like Gorilla and ToolBench that isolate API usage, SOP-Bench evaluates agents on full workflows with interdependencies, ambiguity, and error handling requirements. It provides a more comprehensive testbed for assessing agent robustness in enterprise automation settings.
3 Generating Industry-grade SOPs
At the core of our work lies the SOP-Bench benchmark generation workflow, a novel LLM prompting pipeline that leverages Anthropic’s Claude 3.5 Sonnet v2 model in conjunction with human oversight to develop industrial-grade SOPs across a diverse spectrum of domains that mimics real-world procedures (Figure.  2). The pipeline employs a hierarchical prompting strategy with six specialized templates, that produces SOPs which capture the nuances of human-authored procedures such as industry-specific jargon, interdependent logic, implicit domain knowledge, ambiguous instructions, and cross-referenced content. The framework enables agent builders to systematically create domain specific benchmarks with appropriate datasets and tools needed to evaluate the LLM agent’s performance on sequential reasoning, tool use and constraint adherence in procedural knowledge tasks. The benchmark generation workflow has two major steps: in the first stage, all the components of the benchmark such as SOP, tool specifications and code are generated without introducing any unnecessary challenges, ensuring that the core procedural logic is clearly expressed and coherent. In the second stage, we selectively introduce complexity in order to make the SOP as realistic as possible to real-world conditions by modifying the SOP text to include additional constraints, dependencies, or ambiguities such as incomplete information as well as adding redundant tools to challenge the LLM agent’s reasoning and decision-making in tool selection.
Figure 2: SOP-Bench Synthetic Benchmark Generation Workflow using LLMs and Human oversight. ** denotes post-generation complexity introduction in respective components.
3.1 SOP-Bench Benchmark Generation Workflow
The benchmark generation process uses two inputs from the user namely: <Business Task, Task Context>. These two inputs form the critical backbone of the benchmark generation process. A detailed example is provided in  5
Business Task: A one line description of the business task that clearly articulates the topic and purpose of the SOP to be created.
Task Context: A brief description of the different procedural steps required in the generated SOP in order to process a given user request successfully. For the sample patient intake SOP, this includes steps such as validating patient information, insurance details, previous medical history and assessing risk factors.
Here is an example of business task and task context which includes high-level business objectives that is used to guide SOP generation.
  • Business Task
    • SOP outlining the steps for registering new patients
  • Task Context
    • Describe the purpose and scope of the patient intake process specific to new patients
    • Validate insurance information and eligibility
    • Validate prescription insurance
    • ... Continued ...
3.1.1 Dataset Schema Generation
Instead of directly prompting the LLM with <Business Task, Task Context> to generate SOPs, we first generate a structured dataset schema using one-shot prompting (see Appendix C.2). This schema outlines real-world task scenarios, specifying input parameters, intermediate logic, decision points, compliance checks, and expected outcomes. This step is crucial for minimizing hallucinations and correcting structural inconsistencies, thereby establishing a reliable foundation for SOP and tool generation. Acting as a semantic scaffold, the schema embeds critical operational context that guides and constrains subsequent SOP generation, ensuring outputs are both relevant and executable. With human (domain expert) oversight, we validate that the schema captures essential task components, including implicit knowledge and domain-specific constraints. As illustrated in the following example, When the data schema is not referenced, the SOP instructions remain vague whereas adding the dataset schema ensures instructions are precise and consistent. For example,
  • Schema Entry
    • smoking_statusstring, multiple choice for current smoking habits, Choices = Never, Former, Current
  • SOP Generated Without data schema entry in context
    • ... Calculate the smoking risk index using preferred methodology ...
  • SOP Generated With data schema entry in context
    • ... Calculate the smoking risk index (Never, Former, Current) ...
3.1.2 SOP Document Generation
The Standard Operating Procedure (SOP) document captures all operational steps required to complete a task. Using <Business Task, Task Context, Dataset Schema> (Appendix C.1), we generate a structured SOP that outlines procedure objectives, prerequisites, input requirements, detailed instructions, and decision logic in matural language. The inclusion of task context and a human-verified dataset schema forms a unified semantic scaffold that grounds the SOP in domain-specific structure and preserves causal and operational dependencies. This ensures the resulting SOPs are clear, consistent, and aligned with real-world workflows. Human oversight is used to validate domain standards, task logic, and correct inconsistencies or hallucinations, ensuring the SOPs are both practically executable and contextually appropriate for industrial use.
3.1.3 Dataset Generation
We then utilize <Business Task, Task Context, Dataset Schema, SOP>(Appendix  C.3) to generate the underlying dataset that serves as the foundation for SOP execution and evaluation. The dataset represents full procedural context as well as structured inputs and outputs for every tool/API invocation, task parameters, decision pathways and final outcomes. By generating the dataset using dataset schema and SOP, we ensure that the dataset is contextually grounded. Human oversight is used to ensure each datapoint’s logic is unambiguous, the dataset contains appropriate data-types and accurately reflects the SOP logic. The dataset is constructed to include datapoints that challenge the reasoning, problem-solving, compliance, error-handling and tool selection capabilities of LLM agents by including both positive and negative decision pathways as well as edge conditions and task failure scenarios.
3.1.4 API & Tool Specification Generation
To enable LLM agents to dynamically retrieve task-relevant information, we use the generated datasets to define APIs that map to the operational flow of the SOP. Using <Task Context, Dataset, SOP> (Appendix C.4), we generate APIs specifying required inputs (e.g., patient_name) and expected outputs (e.g., patient_id) (Figure  4 in Appendix B). To inform the agent of available tools at each step, we generate corresponding tool specifications using <Task Context, Dataset, SOP, APIs>, also shown in Figure 4 in Appendix B). Human experts validate that APIs and ToolSpecs align with the dataset schema and access only relevant columns, ensuring semantic consistency and operational validity. Here is an example which demonstrates that without clear specifications, the LLM lacks the necessary context to infer input parameters and expected outputs.
  • API Specification
    • getPatientID – retrieves a unique patient identifier based on a provided name  4
  • Tools Code Generation Without API in context
    • def getPatientID(self):
    • return {}
  • Tools Code Generation With API in context
    • def getPatientID(self, patient_name):
    • # Calls internal hospital system to fetch patient ID
    • return self.patientDB.lookupIDByName(patient_name)
In contrast, explicit API specs provide structure and semantic clarity, enabling the LLM to generate more accurate, robust, and callable tool implementations that integrate correctly into the larger system.
3.1.5 Tools Code Generation
To generate executable code for LLM agents to process tasks, we used <Dataset, APIs> (Appendix C.5) to produce tool code reflecting the logic needed for each procedural step. The dataset snapshot and API specs provide input-output contracts as well as data characteristics, ensuring generated code aligns with operational context and system requirements. By processing all APIs together, the LLM identifies reusable patterns such as validations, conditionals, error handling and distills them into consistent templates to improve code efficiency and reliability.
3.1.6 ** Adding complexity in generated components **
In our benchmark generation workflow, SOP-level complexity is primarily introduced through branching logic (e.g., “If a condition fails, choose path P; else path Q”), requiring AI agents to reason over multiple potential execution paths. Each SOP also consists of a set of tools that the agents must use to execute certain functionalities such as fetching data, verifying status, etc.
To evaluate whether agents can address real-world ambiguities while maintaining high performance when faced with confusing and overlapping instructions in a SOP, we applied a strategy of rewriting a SOP to include additional redundant details that are typically related to the business task, but not directly relevant to executing the core steps of the SOP. These additions mimic real-world documentation noise, partially obfuscating the key steps in the SOP with contextual background, tangential best practices and outdated alternatives, which agents must learn to filter out. In parallel, with some SOPs, we introduced semantically similar but functionally redundant tools, creating scenarios where the agents must critically evaluate each option rather than relying on surface cues.
Together, these interventions, induced either by LLMs via prompting or manual authoring by humans, and validated through human oversight, allow agent builders to rigorously assess their agent’s tools selection behavior, identify failure modes, and iteratively refine both the tool definitions and the agent’s decision-making policy for truly robust SOP comprehension and tool selection.
Table 3: Characteristics of Benchmark SOPs in SOP-Bench. $n$ = number of tasks per SOP; $t$ = number of tools available; $\tau$ = token count, using GPT 3.5 turbo [14]; $\mathbb{C_{H}}$ = human-perceived complexity, average of three domain experts of three dimensions, ease of understanding, ambiguity or implicit domain knowledge and reasoning complexity, on a scale of 1 - 10; $\mathbb{C}_{llm}$ = complexity estimated by Claude 3.5 (scale 1-10, calibrated against Berkeley function calling tasks as baseline 2)
SOP Topic Domain $n$ $t$ $\tau$ $\mathbb{C}_{H}$ $\mathbb{C}_{llm}$
Content Flagging Content Moderation 226 4 850 7.67 9
Customer Service Support 208 10 1457 9 8
Dangerous Goods Supply Chain 327 5 775 5 7
Aircraft Inspection Transportation 150 7 834 6.67 9
Email Intent Retail Seller Mgmt 122 4 766 5.33 7
Know Your Business Finance 122 8 1359 8.33 9
Patient Intake Healthcare 90 6 753 4.33 7
Video Annotation Autonomous Driving 168 26 4492 9.67 10
Video Classification Media 198 25 1249 8.33 9
Warehouse Package Inspection Logistics 200 12 653 4.67 9
Average 181 11 132
Total 1,811 107
3.2 SOP Descriptions
SOP-Bench consists of 10 diverse generic SOPs instantiated in 1800+ tasks that were designed to represent realistic industrial procedures. Each SOP presents unique challenges that test different aspects of agent capabilities, from complex decision-making to tool usage and structured output generation. These SOPs expose agents to a broad spectrum of business modalities that include classification such as categorizing email intent, structured verification based modalities such as business verification, industrial processes such as aircraft pre-flight inspection as well as generic business assignments such as annotation, content flagging, customer service, etc. Each of these SOPs help agent builders to explore different forms of reasoning, compliance-handling, nuanced intent interpretation and procedural task planning and execution. See appendix B for details. The detailed stats are listed in table 3, where we enumerate the topics, domain, task count ($n$), API count ($t$), SOP token count ($\tau$) and rate of complexity levels by we ask three human domain experts to rate the SOPs on (1) ease of understanding (2) ambiguity or implicit domain knowledge required to comprehend the SOP, and (3) reasoning complexity. We average the scores and present in the table. We also ask Claude 3.5 v2 to rate the SOPs on complexity on a scale of 1 - 10.
4 Agents for automating SOPs
To evaluate the effectiveness of LLM-based agents in executing synthetically generated SOPs that mimic their real-world counterparts, we implement two complementary agent architectures in SOP-Bench: (a) Function Calling (FC) Agent and (b) ReAct Agent. Both agents utilize Anthropic’s Claude 3.5 Sonnet v2 model, leveraging Claude’s native tool-calling interface to dynamically invoke external functions/APIs during task execution. More details on these two agents are in Appendix  A.
We evaluate both agents on SOP-Bench using standardized scripts that: (1) load SOP documents and structured task datasets; (2) convert each task into natural language instructions; and (3) execute agents while logging outputs, tool traces, and reasoning steps. To simulate API calls, each dataset includes precomputed inputs and outputs for every tool call, stored as columns. These mocks replace live APIs to enable stable, reproducible evaluation without runtime variability. In deployment, they would be swapped for real APIs. For some SOPs, we also compare intermediate decisions, based on API responses, with expected ground truth to assess multi-step reasoning. Though not returned by tools, these values are logged for evaluating agent reasoning. This setup supports controlled comparisons of agent behavior, tool use, and LLM performance across diverse SOPs.
5 Evaluating Agents with SOP-Bench
We evaluate FC-Agent and ReAct-Agent on SOP-Bench using three inputs: (1) Task, (2) SOP (instructions), and (3) ToolSpecs. The outputs include (1) intermediate tool outputs, (2) final response, and (3) trace of tool use and reasoning. We define: $n$ (total number of tasks per SOP), $n_{c}$ (number of tasks per SOP marked complete by the agent), and $n_{a}$ (number of correctly completed tasks per SOP). Evaluation metrics are: (1) Execution Completion Rate ($ECR$): proportion of tasks the agent marked as complete, $ECR=\frac{n_{c}}{n}$; (2) Conditional Task Success Rate ($C$-$TSR$): fraction of completed tasks that match the ground truth, $C$-$TSR=\frac{n_{a}}{n_{c}}$; and (3) Task Success Rate ($TSR$): overall accuracy, $TSR=\frac{n_{a}}{n_{t}}$. While $ECR$ and $TSR$ serve as core metrics, certain tasks demand domain-specific evaluation. For example, content moderation emphasizes false positive/negative rates, while patient intake prioritizes process adherence and compliance with privacy standards. Though this work reports only $ECR$ and $TSR$, future research should incorporate tailored metrics to better assess agent performance across industry verticals.
Table 4: Agent Evaluation Results on SOP-Bench: ECR = Execution Completion Rate; TSR = Task Success Rate; C-TSR = Conditional Task Success Rate. All percentage values are rounded up to the nearest integer.
SOP Topic $n$ FC-Agent React-Agent
ECR C-TSR TSR ECR C-TSR TSR
Content Flagging 226 0% 0% 0% 90% 11% 10%
Customer Service 208 100% 14% 14% 88% 4% 3%
Dangerous Goods 327 68% 87% 59% 83% 96% 80%
Aircraft Inspection 150 100% 1% 1% 97% 89% 87%
Email Intent 122 78% 39% 31% 60% 100% 60%
Know Your Business 122 100% 19% 19% 92% 39% 36%
Patient Intake 90 33% 0% 0% 100% 0% 0%
Video Annotation 168 100% 49% 49% 70% 99% 69%
Video Classification 198 100% 52% 52% 54% 39% 21%
Warehouse Package Inspection 200 41% 32% 13% 54% 39% 21%
Weighted Average 70% 35% 27% 83% 61% 48%
Figure 3: Shows the relationship between human perceived complexity and task success rates (TSR) of the two agents. There is very low correlation, $\rho(\text{TSR}_{\text{FC}},\mathbb{C_{H}})=0.19$, $\rho(\text{TSR}_{\text{React}},\mathbb{C_{H}})=0.12$, both $p>0.10$, so statistically insignificant, thus highlighting the fact that even SOPs with low human perceived complexity could be complex for LLMs.
Table 4 shows the average  [^3] We have averaged percentages of agents’ success rates over 10 SOPs. With each SOP being different from the others we dont think the error distribution is identical, which violates the normality assumption. Hence, we are not reporting error bars of the average proportion estimates. performance metrics of both FC-Agent and React-Agent on all the tasks in SOP-Bench. On average, ReAct-agent performs better than FC-agent. Such analysis and conclusions are only possible when a benchmark like SOP-Bench is available.
6 Discussion
In explaining our preliminary findings/results, we consider several factors that could affect agent performance, such as (1) number of tools needed, (2) complexity of reasoning required, (3) presence of ambiguous or implicit SOP instructions, (4) need for long-term memory or state tracking, and (5) degree of procedural vs. declarative knowledge involved. For instance, React-Agent performs exceptionally well on tasks like Dangerous Goods Classification SOP and Email Intent SOP, which require nuanced reasoning and decision chaining, but fails completely on tasks like Customer Service and Patient Intake that demand precise form filling and tool invocation—highlighting its difficulty with structured, deterministic workflows. Conversely, the FC-Agent succeeds on tasks where deterministic tool use is sufficient, but struggles when the task requires deeper context inference or adaptive reasoning, as seen in Content Flagging or Know Your Business.
The content flagging SOP which resulted in lowest ECR and TSR, involves four API calls, each requiring five to six input parameters and a strict execution order. The agent must reliably invoke these tools in the correct sequence, accurately populate the input parameters, and retain intermediate outputs in memory for use in subsequent steps. However, analysis of the agent logs reveals frequent failures due to tool invocation errors and inconsistent function outputs. Despite aiming to follow the SOP closely, the agent often deviates—either by skipping steps or manually approximating values when a tool fails. In several instances, the agent resorts to hallucinating results; for example, when the calculate user trust score function fails, it fabricates a value between 0 and 100. These issues lead to lower accuracy (lower TSR).
The evaluation of tool-calling accuracy on two SOPs shows significant variability, highlighting both the challenges and potential of using LLMs for task automation. In the Content Flagging SOP, which uses a smaller set of four tools, the FC-agent successfully invoked all required tools in 66.9% of tasks. However, this includes cases with incorrect parameters, exposing a gap in parameter handling. Despite the high invocation rate, 79.3% of tool calls failed, with 60.6% of failures due to parameter issues like sequence mismatches or invalid argument types. These findings underscore the need for better parameter validation and error handling to enhance reliability.
The Video Classification SOP posed a more complex challenge, with 25 available tools but only 5 required per task. Despite this, the agent invoked at least one incorrect tool in every execution, leading to a 100% rate of unwanted tool calls. Although the agent eventually completed the tasks, frequent irrelevant tool use reduced efficiency and reliability. Overall, 74.8% of tool calls failed, with 50.6% due to incorrect parameters, highlighting the difficulty of parameter alignment in complex, multi-tool settings. As task complexity grows, accurate tool selection and parameter handling become increasingly critical for reliable execution.
Finally, the near-zero TSR for both agents in Patient Intake illustrate another critical challenge: tasks involving sensitive domains (e.g., healthcare) often require strict sequencing, nuanced input parsing, and adherence to implicit compliance policies. Current agents struggle to manage these interleaved constraints without explicit fine-tuning or specialized architectural support.
7 Conclusions and Next Steps
This work presents SOP-Bench, a comprehensive benchmark generation workflow and associated benchmarks that are designed to evaluate AI agents on real-world SOPs across diverse industrial domains. Our preliminary findings with two popular agent architectures validate the central thesis of this work: agent performance is highly context- and task-dependent, and robust benchmarking across realistic SOPs is essential to inform further progress and research.
In a nutshell, SOP-Bench represents a significant leap forward in benchmarking the performance and robustness of AI agents in enterprise automation settings. It enables rigorous evaluation of agents on realistic, industrial-grade SOPs that reflect real-world complexity including task inter-dependencies, procedural ambiguity, conditional logic, and error-handling requirements. Moving forward, there are several promising directions for extending this benchmark:
(1) Introduce multiple versions of the same SOP with varying levels of complexity and abstraction and systematically study how agent performance scales with task difficulty.
(2) Incorporate multiple LLMs in the SOP and data generation pipeline to reduce linguistic and structural bias, and to add richness to procedural variations.
(3) Evaluate additional agent architectures—including retrieval-augmented agents, multi-agent collaborative systems, and memory-augmented agents, to better understand trade-offs and strengths.
(4) Extend SOP-Bench to include multimodal instructions such as images and tables, which are common in enterprise documentation and workflows.
(5) Include hierarchical SOPs — with sections and sub-sections — that involve significant context switching back and forth, thereby adding substantial complexity to the SOPs.
We encourage the research community and practitioners to use SOP-Bench in two primary ways: (1) to evaluate and compare different agent designs for tasks that are structurally or semantically similar to those in their domain, and (2) to adopt our framework for generating task-specific SOP benchmarks when such examples are not readily available in SOP-Bench. In either case, systematic benchmarking should precede production deployment to ensure that the chosen architecture aligns with domain-specific requirements, robustness needs, and compliance constraints.
Broader impact
Positive Societal Impacts
SOP-Bench has the potential to significantly accelerate the deployment of safe and effective LLM agents in high-stakes industrial workflows, where reliability and consistency are paramount. Furthermore, the public release of SOP-Bench, along with the data generation framework, democratizes access to high-quality benchmarks for both academic and industry researchers. This can foster transparent development, reproducible evaluation, and community-driven progress toward safer AI systems for enterprise applications.
Negative Societal Impacts
There is a risk that widespread adoption of SOP-following agents may lead to over-reliance on automated decision-making in sensitive domains, such as healthcare or legal triage, without adequate human oversight. If not rigorously tested for edge cases and bias, LLM agents could propagate errors, reinforce institutional inequities, or introduce new failure modes in mission-critical systems.
We encourage responsible use of SOP-Bench, especially in regulated or high-risk domains, and advocate for human-in-the-loop solutions and evaluations in all downstream applications.
Disclaimer
IMPORTANT: The Standard Operating Procedures (SOPs) used in this benchmark are synthetically generated using Large Language Models for research and evaluation purposes only. These SOPs are designed to test AI agent capabilities and DO NOT represent actual operational procedures from any real organization. They should NOT be used or adapted for real-world implementation or deployment. The authors make no claims about the safety, completeness, or regulatory compliance of these synthetic SOPs. Similarly, all example data, including but not limited to patient records, user profiles, transaction logs, and any other structured or unstructured data used in this benchmark, are entirely synthetic and AI-generated. This data does not represent or correspond to any real individuals, organizations, or actual scenarios. Any resemblance to real persons, or organizations, or situations is purely coincidental. This synthetic data is provided solely for research purposes and should not be used as a basis for real-world decision-making or system implementation.
Appendix
Appendix A Technical Details on Agents
A.1 FC Agent
This is a prompt-driven, lightweight execution engine leveraging the native tool calling capability of LLM (e.g, Claude). It accepts an SOP text and a structured task input, dynamically generates prompts using a parameterized template engine, and maintains a conversation loop with Claude throughout the execution. During agent initialization, it registers external tools defined in an OpenAPI-style tools specifications JSON file. During execution, when the model triggers a tool_use event, the agent extracts the relevant tool name and parameters, executes the corresponding function via a domain-specific ToolsManager, and delivers the output to the model through a toolResult message. A built-in self-reflection mechanism automatically responds to non-critical clarification requests from the model, thereby reducing the need for manual intervention. Final outputs are enclosed in XML-style tags (e.g., <final_response>) and are parsed into structured formats for downstream evaluation. The full FC-Agent prompt template is included in Appendix  D.1 (Listing  6). This architecture emphasizes traceability, minimal supervision, and broad applicability across various SOP domains.
A.2 ReACT Agent
Built on the LangChain ReAct framework [30], this agent is designed to handle SOPs involving complex branching logic, intermediate reasoning, and decision pathways. It ingests the SOP text, structured task input, and tool specifications JSON, similar to the FC Agent. Tools are dynamically discovered and wrapped using @tool annotation to ensure compatibility with the LangChain agent execution framework. The ReAct Agent interleaves reasoning traces (Thoughts) with explicit tool calls (Actions), which are executed via LLM’s tool-calling interface. The resulting Observations are fed back into the reasoning loop, enabling iterative planning and dynamic adaptation to SOP-specific control flows. This process continues until a conclusive Final Answer is reached, marked using tags such as <final_decision> and <final_decision_reason>. The system and user prompt templates used to steer the ReAct Agent are detailed in Appendix  D.2 (Listings  7 and  8). This agent architecture supports more sophisticated workflows, enabling adaptive behavior in SOPs requiring fallback logic, verification steps, or multi-phase reasoning.
Appendix B Details of the SOPs in SOP-Bench
In this section, we briefly describe the diverse industry use cases for which we generated synthetic data, and drawing parallels to real-world operational workflows.
B.1 Content Flagging
This SOP models a content moderation workflow designed to evaluate flagged user-generated content through a combination of bot detection, trust scoring and violation severity assessment. The SOP mimics a workflow in use cases such as fraud detection in fintech or abuse prevention in social platforms, where automated AI and software systems have scored the content and the results of the upstream system are available as input to the human agent. The SOP integrates behavioral heuristics (e.g., captcha interaction, device fingerprinting), geolocation risk analysis, and multi-dimensional scoring frameworks such as the Bot Probability Index (BPI), User Trust Score, and Content Severity Index, etc. The process culminates in a final decision—ranging from content removal to user banning—based on a rule-based decision matrix. Its complexity arises from the interplay of user behavior history, system signals, and threshold-based branching logic, reflecting the nuanced reasoning required in high-stakes moderation pipelines.
This SOP is highly relevant because it simulates a structured yet non-trivial decision-making process involving conditional logic, multi-step tool usage, and real-world signal fusion. It relies on four distinct APIs (tools) that mirror common platform trust and safety operations, such as bot detection, trust scoring, severity assessment, and final content dispositioning. The evaluation criteria require agents to correctly compute decision outcomes and generate structured audit trails, offering a clear and rigorous way to assess agent planning accuracy, execution fidelity, and interpretability.
B.2 Customer Service
This SOP defines an offline workflow for diagnosing and resolving customer-reported service issues based solely on initial inputs—such as an account ID and issue description—without real-time customer interaction. It is designed for backend support scenarios where resolution relies entirely on internal system data, logs, and historical telemetry, as commonly seen in telecommunications and infrastructure-based services. The procedure covers authentication validation, account eligibility checks, outage detection, diagnostic testing, troubleshooting, and escalation, using predefined thresholds for key metrics (e.g., latency, jitter, bandwidth). Traceability is ensured through session tokens, ticket IDs, timestamps, and audit logs.
This SOP is particularly suited for evaluating AI or human agents in semi-autonomous support systems, where success depends on accurate decision-making, conditional branching, and proper documentation. It simulates a realistic diagnostic environment through a unified API-like framework. Agents must interpret system signals (e.g., suspension status, outage presence, signal quality) and act accordingly. The final output is a structured JSON-based Resolution Summary Document (RSD), encoding the complete decision trail for reproducibility, auditability, and downstream integration.
B.3 Dangerous Goods Classification
This benchmark includes a sophisticated SOP for dangerous goods classification in supply chain operations. The procedure integrates multiple data sources (Safety Data Sheets, Handling Guidelines, Transportation Requirements, and Disposal Guidelines) through a structured scoring system, where each component contributes a severity score (1-5) to determine the final hazard classification. The SOP incorporates practical considerations such as missing data handling, validation checks, and API integrations, culminating in a final classification system (Classes A through D) based on cumulative hazard scores. This example demonstrates the intricate decision-making processes and interdependencies typical in industrial settings, requiring AI agents to handle document parsing, numerical computations, conditional logic, and structured output generation - capabilities that go far beyond simple task execution.
Similar workflow patterns and complexity are found across various industrial applications, such as pharmaceutical quality control processes, aircraft pre-flight safety checks, food safety certifications, and industrial equipment maintenance prioritization. These use-cases share key characteristics including multi-source data integration, quantitative scoring systems, complex decision trees, comprehensive documentation needs, and API integrations, demonstrating how our benchmark represents realistic challenges faced in industrial automation.
B.4 Video Annotation for self driving
As we rapidly move to a world powered by AI where autonomous vehicles navigate our streets, there is a need to produce high-quality datasets to train perception and navigation systems. This serves as the main motivation behind creating a SOP that outlines steps to detect objects and annotate videos. The SOP establishes input requirements (format, frame rate, resolution, color depth), comprehensive processes for object detection and segmentation using specialized tools, and strict quality control measures (spatial accuracy, temporal consistency, inter-annotator agreement). After processing, the agent must report whether the video was successfully processed, provide paths to the resulting object detection and segmentation files, and include human reviewer scores to validate annotation quality.
B.5 Know Your Business
Verification of business entities is an essential component in the modern economy. This serves as the primary motivation for creating a comprehensive SOP focused on business verification that identifies high-risk entities before engaging in business relationships. The SOP outlines step-by-step guidelines on how to validate business details, registration, licenses, and tax identities, as well as ownership structures and sanctioned individuals. The SOP requires the agent to reason about completeness of business details and call relevant tools to fetch required data to aid in the verification process. Finally, the agent is required to either approve the business or escalate the case for human intervention.
B.6 Video Flagging for Content Moderation
In the online world of content consumption and social media, content moderation plays a critical role in ensuring user safety, maintaining community standards, and preventing the spread of harmful or misleading information. As platforms scale, the volume and variety of user-generated content increase exponentially, making manual moderation infeasible. This has led to the growing reliance on automated moderation systems powered by artificial intelligence and machine learning. This SOP establishes a comprehensive framework for the systematic classification, escalation, and moderation of user-generated video content on digital platforms. The Agent uses tools to fetch initial review metadata and uses reasoning to escalate the content to human moderator and based on their feedback, assigns moderation tags, content warning status and a final decision to ultimately keep or take down the video.
B.7 Patient Intake
This synthetic SOP models the end-to-end intake and registration process for new patients in clinical settings, incorporating insurance validation, prescription benefits verification, lifestyle and clinical risk stratification, and pharmacy network checks. Its complexity lies in the conditional decision pathways, cross-system interactions (e.g., mock prescription and pharmacy validations), and the integration of structured data inputs with algorithmic assessments like risk scoring. The SOP requires the agent to perform validation tasks using six APIs—each mapping to a distinct functional step—and to generate structured outputs that reflect real-world registration outcomes under healthcare compliance constraints.
As a benchmark dataset for evaluating LLM-based task planning agents, this SOP is highly relevant due to its realistic structure and deterministic yet context-sensitive branching logic. It simulates challenges found in broader industrial domains such as healthcare operations, insurance underwriting, and benefits management. Similar real-world workflows include hospital patient onboarding, insurance claims adjudication, and pharmacy eligibility reviews. The SOP’s evaluation framework scores agents on whether they produce complete, correctly formatted outputs across six output fields and the final decision on patient registration, aligning well with goals of assessing LLM agent task planning, reliability, and tool execution precision.
Figure 4: Sample API and ToolSpec generated for the SOP related to Patient Intake in Healthcare industry
Figure 5: Detailed Sample of Business Task and Task Context for the Patient Intake SOP relevant to healthcare industry
B.8 Seller email intent
This SOP is designed for processing seller appeals in e-commerce platforms, representing the complexity of modern customer communication systems. The procedure systematically handles email classification and response determination through a structured workflow that integrates multiple data sources and API endpoints. The SOP instructs categorizing seller communications into distinct categories (pricing concerns, description modifications, listing status inquiries, or general questions) and validate these against product databases, pricing systems, and inventory management through specific API calls. The procedure incorporates practical considerations such as data validation, error handling, and compliance documentation, culminating in standardized XML-formatted outputs with clear action items. This example demonstrates the intricate decision-making processes typical in modern business operations, requiring AI agents to handle NLU, multi-system data integration, and protocol-based response generation. Similar workflow patterns and complexity are found across various industries, such as customer support ticket systems, insurance claim processing, financial services communication handling, and travel booking support systems. These use-cases share key characteristics including intent classification needs, multiple API integrations, structured decision trees, standardized response protocols, and compliance documentation requirements.
B.9 Warehouse Package Inspection
This SOP enhances the warehouse package inspection workflow by introducing a structured protocol for detecting and resolving shipment discrepancies. It reflects real-world operational needs in enterprise logistics, where issues like barcode mismatches, quantity variances, and damaged goods require swift, automated handling to maintain inventory integrity and vendor accountability.
The process begins with barcode validation using image comparison. If the barcode does not match the confirmed product ID, the item is immediately classified as “Wrong Item” and marked for return, bypassing further checks. Otherwise, the agent calculates quantity variance and classifies discrepancies based on predefined rules—such as overage, underage, cancelled quantity, or severe mismatch when the variance exceeds a set threshold.
Subsequent checks include verifying warehouse location and assessing package condition via image-based damage detection. Discrepancies like misdirected shipments or visible damage are accordingly classified. Once all issues are identified, the system calculates financial penalties using a chargeback matrix, applying formulas based on discrepancy type and unit cost.
The procedure concludes with structured outputs: a problem classification report, chargeback sheet, and resolution status update. Designed to simulate operational flows in high-volume retail logistics, this SOP challenges agents to integrate vision processing, rule-based logic, and structured reporting—providing a rigorous benchmark for LLM agents in industrial SOP execution.
B.10 Aircraft Inspection
This SOP establishes a framework for pre-flight airworthiness verification of aircrafts, covering multi-layered inspections like mechanical component checks, electrical system authentication, and maintenance record validation through seven distinct API tools. This SOP is valuable for evaluating LLM-powered agents, as it tests their ability to interpret technical parameters, manage complex multi-step workflows, and produce structured outputs for audit purposes. The complexity arises from its complex input structure and conditional branching (e.g., reporting inspection mismatches where found). It also has cross-domain applicability in industries like automotive manufacturing, healthcare equipment maintenance, and industrial machinery inspection, where precise validation and discrepancy reporting are critical.
Appendix C Prompt Templates
C.1 SOP Document Generation Prompt
Listing 1: Prompt Template YAML for converting user-provided task title and context to SOP
1prompts:
2 - name: complex_sop_generator
3 parameters: [sop_title, additional_context, sample_schema]
4 text: |
5 You are an expert AI system specialized in creating highly detailed Standard Operating Procedures (SOPs) for complex, technical, and regulated environments. Your task is to generate an SOP that is rich in domain-specific terminology, intricate logic, and advanced structure-requiring substantial subject-matter expertise to fully comprehend.
6
7 Here is the title for the SOP you need to create:
8 <sop_title>
9 {{sop_title}}
10 </sop_title>
11
12 {if additional_context != "" %}
13 Here is additional context that you must incorporate into the SOP generation process:
14 <additional_context>
15 {{additional_context}}
16 </additional_context>
17 {endif %}
18
19 {if sample_schema != "" %}
20 Here is a schema for the SOP that you should use to guide your creation process:
21 <sample_schema>
22 {{sample_schema}}
23 </sample_schema>
24 {endif %}
25
26 Before generating the SOP, you must thoroughly analyze the title, additional context and sample schema to plan the most sophisticated and complex approach possible. Conduct this analysis and planning phase inside <sop_analysis> tags. Follow these steps:
27
28 1. Deconstruct the title and additional context into key components.
29 2. Identify the specific industry or field the SOP pertains to.
30 3. Enumerate potential regulatory requirements relevant to this SOP.
31 4. List out key technical terms that will be used in the SOP.
32 5. Brainstorm complex technical details for each section of the SOP.
33 6. Outline interdependencies between different sections of the SOP.
34 7. Consider potential complications, edge cases, and rare scenarios that should be addressed.
35 8. Identify specific compliance issues and how they will be addressed in the SOP.
36
37 Use industry-specific jargon and complex terminology throughout the planning process. Consider industry-specific nuances, potential regulatory requirements, and intricate technical details that could be incorporated. This planning phase should be extensive and detailed.
38
39 After completing the planning phase, generate the SOP using the following structure:
40
41 <SOP>
42 1. Purpose
43 [Provide a clear, technical statement of the purpose of SOP]
44
45 2. Scope
46 [Define the exact coverage of the SOP and specify to whom it applies]
47
48 3. Definitions
49 [List and define relevant terms, ensuring they are technical and field-specific]
50
51 4. Input
52 [Enumerate required inputs, materials, or information needed to initiate the procedure]
53
54 5. Main Procedure
55 5.1 [First main step]
56 5.1.1 [Substep with detailed explanation]
57 5.1.2 [Substep with detailed explanation]
58 5.2 [Second main step]
59 5.2.1 [Substep with detailed explanation]
60 5.2.2 [Substep with detailed explanation]
61 [Continue with all necessary steps and substeps, ensuring each is explained in detail]
62
63 6. Output
64 [Describe the expected results or deliverables upon completion of the procedure]
65 </SOP>
66
67 Requirements for SOP Generation:
68
69 1. Use advanced industry jargon, long-form technical sentences, and assume your audience has deep domain expertise.
70 2. Ensure that the language used throughout the SOP is sophisticated and technical, requiring a high level of expertise to fully comprehend.
71 3. Include cross-references between different sections to create a web of interconnected information.
72 4. In the Definitions section, include complex terms that will be used throughout the SOP, ensuring they are technical and field-specific.
73 5. Use the information provided in the additional context to construct the SOP.
74 6. Do not use bullet points or lists. Instead, express each item in fully elaborated prose, forming complete paragraphs. Each requirement or specification must be described with technical detail, rationale, and interconnections, mimicking the tone of formal documentation or academic writing.
75 7. At each stage, explicitly specify quantitative thresholds and conditions that guide decisions (e.g., scoring cutoffs, routing rules, escalation triggers).
76 8. The SOP should be self-sustaining, all necessary instructions to complete the task must be explicitly mentioned in the SOP
77 9. If a sample schema is provided, use the information exhaustively to design thresholds, success and failure logic to guide decision-making in the SOP, increasing its complexity.
78 10. For each section, use the <sop_analysis> tags to think and reason about ways to increase complexity and depth. Consider potential pitfalls, rare scenarios, and intricate details that might not be immediately obvious.
79 11. Generate the entire SOP in one continuous pass without pausing, asking for confirmation, or requesting interaction. Do not ask if you should continue. Output the full SOP all at once within the <SOP> </SOP> tags.
80
81 Remember to make each step and substep in the Main Procedure detailed and clearly explained, while maintaining the required level of complexity and technical depth. This is crucial for the effectiveness of the SOP in its intended environment.
C.2 Data Schema Generation Prompt
Listing 2: Prompt Template YAML for converting user-provided task title, task context and a one-shot demonstration of a dataset schema to Synthetic Dataset Schema
1prompts:
2 - name: generate_dataset_schema
3 parameters: [sop_title, additional_context]
4 text: |
5 You are an expert AI assistant that specializes in generating synthetic dataset schemas based on Standard Operating Procedures (SOPs).
6 You will use the Standard operating procedure(SOP) text to generate a dataset schema.
7
8 First, carefully read through the following SOP title:
9
10 <sop_title>
11 {{sop_title}}
12 </sop_title>
13
14 {if additional_context != "" %}
15 Here is additional context you must adhere to for the SOP generation process:
16 <additional_context>
17 {{additional_context}}
18 </additional_context>
19 {endif %}
20
21 You must structure the output using this format:
22 <schema>
23 video_id - string, unique identifier for the video (e.g., "vid_00123").
24 video_path - string, file path or URI of the video (e.g., "/data/videos/vid_00123.mp4").
25 upload_timestamp - datetime, when the video was uploaded (e.g., "2025-04-18T15:32:10Z").
26 uploader_id - string, anonymized user identifier (e.g., "user_487").
27 video_language - string, detected or declared language of spoken/audio content (e.g., "en", "es").
28 region - string, region associated with the video or uploader (e.g., "US", "IN").
29 metadata_tags - list[string], keywords or tags extracted from video metadata (e.g., ["protest", "crowd", "chanting"]).
30 format_validated - boolean, whether the video passed format checks (e.g., True).
31 duration_seconds - integer, total length of the video in seconds (e.g., 312).
32 </schema>
33
34
35 Now, Generate a dataset schema within <schema></schema> tags for the given SOP title using the additional context .
36 <requirements>
37 1. Define the schema of the dataset, including each field:
38 - Name
39 - Type (e.g., string, integer, float, boolean, datetime, enum, list, etc.)
40 - Description (what the field represents)
41 - Example value
42 2. Include data ranges and choices wherever applicable.
43 </requirements>
C.3 Dataset Generation Prompt
Listing 3: Prompt Template YAML for converting user-provided task title, task context and dataset schema to Synthetic Dataset
1prompts:
2 - name: generate_dataset_csv
3 parameters: [sop_title, sop_file_contents, additional_context, sample_schema, n_samples, sop_data_generation_guidelines]
4 text: |
5 You are an expert AI assistant trained to generate high-quality synthetic datasets based on detailed Standard Operating Procedures (SOPs). Your task is to extract operational logic and decision conditions from the SOP to synthesize structured data that reflects the full range of possible cases.
6 You will be provided with the following inputs:
7
8 <sop_title>
9 {{sop_title}}
10 </sop_title>
11
12 <sop_contents>
13 {{sop_file_contents}}
14 </sop_contents>
15
16 {if additional_context != "" %}
17 <additional_context>
18 {{additional_context}}
19 </additional_context>
20 {endif %}
21
22 {if sample_schema != "" %}
23 <schema>
24 {{sample_schema}}
25 </schema>
26 {endif %}
27
28 Your task is to generate a synthetic dataset as follows:
29
30 1. Interpretation: Carefully analyze the SOP to identify all decision points, validation steps, escalation conditions, and classifications. Use these to determine what fields must appear in the dataset and what values those fields may take.
31 2. Schema-Adherence: Ensure the dataset reflects and expands upon the fields in the ‘<schema>‘ section, if provided. If no schema is provided, infer a realistic and complete schema from the SOP contents.
32 3. Diversity of Cases:
33 - Include both valid and invalid examples.
34 - Cover edge cases, borderline values, and policy violation cases.
35 - Ensure each step of the SOP’s logic is reflected in at least some rows.
36 4. Data Format:
37 - Output the dataset within ‘<data></data>‘ tags only.
38 - Use Python list-of-lists format including the header row, suitable for immediate use with CSV or Pandas.
39 - Avoid any extra commentary, formatting, or explanations outside the ‘<data>‘ block.
40 5. Validation Requirements:
41 - All rows must be non-empty.
42 - All columns must be complete and contain valid, well-formed values.
43 - No unterminated strings, no newlines within strings.
44 - Ensure consistency between field values (e.g., escalation should only be True when required conditions are met).
45 6. Generate at least ‘{{n_samples}}‘ valid rows in total, spanning the full spectrum of SOP-defined conditions.
46 7. Generate the dataset in one continuous pass without pausing, or asking for confirmation, or requesting interaction. Do not ask if you should continue.
47 8. Do not include any explanatory notes, placeholder comments, ellipses, or messages like "[continuing...]" or "[...more rows below...]". Output only the full dataset rows within the ‘<data></data>‘ block without interruption. Any such messages will be treated as format violations.
48 9. You must ensure that the data generated does not suffer from malformed node or string errors.
49 10. Make sure there are no empty data columns or rows. Make sure there are NO unterminated string literals. Do not add new lines.
50 11. For columns containing boolean, you must ensure that to use True or False so that it can be parsed using ast
51 12. You must ensure you close any square brackets.
52
53 {if sop_data_generation_guidelines != "" %}
54 Additional requirements include:
55 {{sop_data_generation_guidelines}}
56 {endif %}
57
58 <data>
59 [["field_1", "field_2", ..., "field_N"],
60 ["value_1_1", "value_1_2", ..., "value_1_N"],
61 ...
62 ["value_k_1", "value_k_2", ..., "value_k_N"]]
63 </data>’
C.4 SOP API Generation Prompt
Listing 4: Prompt Template YAML for converting user-provided task title, task context and sample data to APIs and Toolspec JSON
1prompts:
2 - name: sop_api_generator
3 parameters: [sop_file_contents, additional_context, sample_data]
4 text: |
5 You are an expert AI assistant specializing in API documentation and generation. Your task is to create sample API calls based on a Standard Operating Procedure (SOP) document. The goal is to produce accurate and comprehensive API documentation that captures all necessary steps and dependencies in the SOP.
6
7 First, carefully read through the following SOP contents:
8
9 <sop_contents>
10 {{sop_file_contents}}
11 </sop_contents>
12
13 {if additional_context != "" %}
14 Here is additional context you must adhere to for the SOP generation process:
15 <additional_context>
16 {{additional_context}}
17 </additional_context>
18 {endif %}
19
20 {if sample_data != "" %}
21 Here is a sample data for the SOP:
22 <sample_data>
23 {{sample_data}}
24 </sample_data>
25 {endif %}
26 Now, conduct a thorough analysis of the SOP and additional context, focusing on the ’Main Procedure’ section, and its sub-sections as well as the workflow. Wrap your analysis in <sop_breakdown> tags:
27
28 <sop_breakdown>
29 1. List all distinct tasks or operations in the ’Main Procedure’ section.
30 2. For each task, identify any external dependencies required for execution.
31 3. Consider and note potential edge cases for each task.
32 4. From the SOP, identify and list the key API parameters for each task.
33 5. Plan your approach for API generation based on this analysis.
34 6. Identify potential sources for input parameters (either from the SOP Input or from other API responses).
35 7. Consider how different APIs might depend on each other and in what order they should be executed.
36 8. For each potential API, list possible error scenarios and how they might be handled.
37 9. from the sample data, list all the available columns.
38 </sop_breakdown>
39
40 <requirements>
41 After completing your analysis, generate a set of tools/APIs based on the SOP. Follow these guidelines:
42 1. Create a separate API for each distinct task or operation in the procedure.
43 2. Ensure each tool or API is as granular or atomic as possible.
44 3. Do not repeat tools or APIs for the same task.
45 4. Capture all possible external dependencies the SOP needs to execute successfully.
46 5. Ensure that input request parameters for each API are derived either from the SOP Input or from the response parameters (output) of other API calls.
47 6. Make sure the APIs are complete and fully functional once implemented.
48
49 For each API you generate, provide the following details:
50 1. Name of the API
51 2. API description
52 3. Endpoint URI
53 4. HTTP Method
54 5. Request Body (including all necessary parameters) that is present in sample data within <sample_data> tags.
55 6. Response (look for corresponding status columns for each API) that is present in sample data within <sample_data> tags.
56 7 Do not add any parameters in Request and Response Body that is not present in sample data within <sample_data> tags.
57 8. Dependencies (if this API depends on the output of other APIs, list them here)
58 9. Potential error scenarios and suggested error handling
59 </requirements>
60
61 Wrap each tool/API inside <API></API> tags, and enclose the entire set of tools/APIs within <TOOLS></TOOLS> tags.
62
63 Here is an example of the expected output structure:
64
65 <TOOLS>
66 <API>
67 Name: [API Name]
68 Description: [Brief description of the API function]
69 Endpoint: [URI]
70 Method: [HTTP Method]
71 Request Body:
72 {
73 [JSON structure of request parameters]
74 }
75 Response:
76 {
77 [JSON structure of response data]
78 }
79 Dependencies: [List of dependent APIs, if any]
80 Error Scenarios:
81 - [Potential error 1]: [Suggested handling]
82 - [Potential error 2]: [Suggested handling]
83 </API>
84 </TOOLS>
85
86 It is OK for the <sop_breakdown> and <TOOLS> sections to be quite long, as they need to thoroughly cover all aspects of the SOP and resulting APIs.
87
88
89
90 - name: generate_json_schema_for_each_api
91 parameters: [sop_file_contents, additional_context, each_api_metadata, sample_data]
92 text: |
93 You are an expert AI assistant tasked with generating a JSON Schema for an API tool based on provided metadata. This tool is one of the many tools that will be used to execute a Standard Operating Procedure (SOP). Your goal is to create an accurate and detailed schema for this tool that aligns with the SOP and meets the specified requirements.
94
95 First, please review the contents of the SOP:
96
97 <sop_contents>
98 {{sop_file_contents}}
99 </sop_contents>
100
101 {if additional_context != "" %}
102 Here is additional context you must adhere to for the SOP generation process:
103 <additional_context>
104 {{additional_context}}
105 </additional_context>
106 {endif %}
107
108 {if sample_data != "" %}
109 Here is a sample data for the SOP:
110 <sample_data>
111 {{sample_data}}
112 </sample_data>
113 {endif %}
114
115 Now, you will be provided with the metadata for a single API tool. Your task is to generate a JSON Schema for this tool. Here’s how you should proceed:
116
117 1. Carefully read and analyze the API metadata.
118 2. Generate the following for the API tool:
119 a. name: A concise, descriptive name for the tool.
120 b. description: A clear explanation of the tool’s purpose and functionality.
121 c. inputSchema: A JSON schema format specifying the tool’s input’, including type’, properties’, and required fields, all under the key json’.
122 3. Ensure that the schema aligns with the overall SOP direction and any specified restrictions.
123
124 Here is the API metadata you will be working with:
125
126 <api_metadata>
127 {{each_api_metadata}}
128 </api_metadata>
129
130 Before generating the final JSON Schema, wrap your analysis inside <sop_tool_analysis> tags:
131 1. Quote relevant parts of the SOP that relate to the tool purpose.
132 2. List potential input parameters based on the SOP and metadata.
133 3. Consider any constraints or special considerations mentioned in the SOP.
134 4. Outline how the tool aligns with the overall SOP direction.
135 5. Identify potential edge cases or limitations of the tool.
136 6. Describe how this tool fits into the larger workflow outlined in the SOP.
137
138 After your analysis, generate the complete JSON Schema for the API tool. The schema should be formatted as an object with a "toolSpec" property, containing "name", "description", and "inputSchema" as shown in the example below:
139
140 {
141 "toolSpec": {
142 "name": "example_tool_name",
143 "description": "A clear description of the tool’s purpose and functionality.",
144 "inputSchema": {
145 "json": {
146 "type": "object",
147 "properties": {
148 "parameter_name": {
149 "type": "string",
150 "description": "A detailed description of this parameter."
151 }
152 },
153 "required": ["parameter_name"]
154 }
155 }
156 }
157 }
158
159 Remember to:
160 - Ensure all required fields are listed in the "required" array within the input schema.
161 - Provide detailed descriptions for each property in the input schema.
162 - Use appropriate data types (string, integer, etc.) for each property.
163 - Maintain consistency in formatting and structure.
164
165 Present your final JSON Schema inside <SCHEMA></SCHEMA> tags."
C.5 SOP API Function Code Prompt
Listing 5: Prompt Template YAML for implementing the functional code for an API, given sample data and a one-shot demonstration of an API along with its implementation and sample data as input
1prompts:
2 - name: llm_coder
3 parameters: [api, sample_data, api_example, coding_example, data_example]
4 text: |
5 You are an expert AI coding system who is capable of writing code in python using popular libraries.
6 You are given the metadata of an API and a sample dataset.
7 Your task is to write a Python function that simulates the APIs.
8
9 Here is an example of an API, dataset and its corresponding implementation.
10 <example>
11 Here is the api metedata:
12 {{api_example}}
13
14 Here is the dataset head in tsv:
15 {{data_example}}
16
17 Here is the corresponding code:
18 {{coding_example}}
[…]
20
21 <requirements>
22 Requirement 1: Generate relevant Python code within <code></code> tags after closely studying the example provided in <example> tags.
23 Requirement 2: You must write test cases as demonstrated in the example.
24 Requirement 3: Import all required Python libraries. Use descriptive and meaningful variable names.
25 Requirement 4: Name the function exactly as specified in the API specifications.
26 Requirement 5: Conform strictly to the API definitions - the fields in the request must map to the function arguments, and the fields in the response must be returned.
27 Requirement 6: Use only the existing columns present in the dataset.
28 Requirement 7: Load the CSV dataset into memory before accessing the required fields.
29 Requirement 8: Implement the process_tool_call() function exactly as demonstrated in the <example> tags.
30 Requirement 9: Generate the code in one continuous pass without pausing, asking for confirmation, or requesting interaction. Do not ask if you should continue.
31 </requirements>
32
33 Following the above example, write code for the following API and dataset to complete the task successfully.
34 <api>
35 API metadata:
36 {{api}}
37 </api>
38
39 <sample_data>
40 Here is the pandas DataFrame head (5 rows):
41 {{sample_data}}
42 </sample_data>
Appendix D Agent Prompt Templates
D.1 Function Calling Agent (FC-Agent)
Listing 6: Prompt Template YAML for FC-Agent to process input request following an SOP
1prompts:
2 - name: simple_sop_agent_v1
3 parameters: [SOP, USER_REQUEST]
4 text: |
5 You are an AI assistant responsible for processing user requests according to a specific Standard Operating Procedure (SOP). Your primary goal is to follow the SOP meticulously, ensuring consistency and accuracy in your responses.
6
7 First, carefully read and internalize the following SOP:
8
9 <sop>
10 {{SOP}}
11 </sop>
12
13 Now, you will receive a user request to process. Your task is to handle this request by strictly adhering to the SOP provided above. Follow these steps:
14
15 1. Read the user request carefully:
16 <user_request>
17 {{USER_REQUEST}}
18 </user_request>
19
20 2. Before responding, review the SOP once more to ensure you understand all steps. In <sop_analysis> tags:
21 a. List the key points of the SOP.
22 b. Identify any potential challenges or edge cases in applying the SOP to the user request.
23 c. Consider possible interpretations of any ambiguous parts of the request.
24
25 3. **CRITICAL:** Process the user request by following each step of the SOP in order. DO NOT skip any steps or add extra steps that are not in the SOP.
26
27 4. For each step, perform the following actions:
28 a. Log your progress in a <step_progress> tag, including the step number, description, and outcome (completed, failed, or pending).
29 b. For critical steps, wrap your analysis inside <analysis> tags to show your calculations, validations, and reasoning.
30 c. If a critical step cannot be completed, state the issue clearly and halt the process before proceeding to the next step.
31
32 5. After every 3 steps, provide a checkpoint summary in <analysis> tags, highlighting critical validations performed and any discrepancies or issues encountered.
33
34 6. If at any point you’re unsure about how to proceed or if the user request doesn’t align with the SOP, state this clearly in your response and halt the process.
35
36 7. Once all steps are completed successfully, provide a final summary of the processed request in <final_response> tags, including all the key information as per the SOP.
37
38 Example output structure:
39
40 <step_progress>
41 Step 1: [Description] - Outcome: [Completed/Failed/Pending]
42 </step_progress>
43
44 <analysis>
45 [Detailed calculations, validations, or reasoning for critical steps]
46 [Checkpoint summaries after every 3 steps]
47 </analysis>
48
49 <final_response>
50 [Final response based on SOP processing]
51 </final_response>
52
53 Remember to adhere strictly to the SOP and provide clear, detailed explanations for each step of the process. It’s OK for the <sop_analysis> section to be quite long.’
D.2 Prompt Template for React Agent
Listing 7: System Prompt Template for React-Agent to process input request following an SOP
1<system_prompt>You are a helpful assistant that follows Standard Operating Procedures (SOPs) to solve problems.
2
3You will be given an SOP in an SOP block <sop> and a list of tools in a <tools> block.
4
5Here are the tools:
[…]
7<tools>
8{tools}
9</tools>
10
11<tool_input_guideline>
12Follow these guidelines when using tools:
131. Examine the tool descriptions carefully to understand what each tool does
142. For each tool, provide all parameters in the correct order and format
153. Format tool inputs as valid JSON objects with all required fields
164. For tools that require nested objects, format them properly as JSON objects
175. Always check that your input includes all required parameters before using a tool
18</tool_input_guideline>
19
20Use the following format:
21<format>
[…]
27</format>
28
[…]
30eventually, once there are no more steps to execute or if the answer cannot be found based on the provided sop, tools and input.
31
[…]
34<final_decision>FINAL_DECISION</final_decision>
35<final_decision_reason>reason for final decision</final_decision_reason>
36</system_prompt>
37
38Begin:
[…]
40{input}
41
42{agent_scratchpad}
Listing 8: A sample User Prompt Template for React-Agent to process input request following an SOP
[…]
Appendix E Examples of Real SOPs
Figure 6: Examples of patient registration SOPs
Appendix F SOP-Bench: Patient Intake and Registration
This appendix provides an example of a Standard Operating Procedure (SOP) package for patient registration in a healthcare setting. It includes the main SOP document, JSON tool specifications, and sample patient data, illustrating how operational procedures are formalized and implemented in healthcare information systems.
F.1 SOP: Patient Intake and Registration
Patient Intake and Registration SOP - Part 1
1. Purpose
This Standard Operating Procedure establishes a framework for intake
and registration of new patients, encompassing demographic data
collection, insurance verification, risk stratification, and pharmacy
network validation in compliance with HIPAA regulations.
2. Scope
This procedure applies to all healthcare personnel involved in patient
registration, including Patient Access Representatives, Insurance
Verification Specialists, and Clinical Intake Coordinators.
3. Definitions
3.1 Protected Health Information (PHI): Individually identifiable
    health information as defined by HIPAA
3.2 Benefits Coordination Protocol (BCP): Systematic verification of
    primary and secondary insurance benefits
3.3 Clinical Risk Index (CRI): Composite score derived from medical
    history and lifestyle factors
3.4 Pharmacy Benefit Management (PBM) Validation: Real-time
    verification of prescription coverage
3.5 Network Adequacy Verification (NAV): Assessment of provider
    network status
3.6 Risk Stratification Algorithm (RSA): Mathematical model for
    patient risk assessment
4. Input
4.1 Required Documentation (some are optional)
- Government-issued photo identification (optional)
- Current insurance card(s)
- Complete medical history questionnaire
- Pharmacy benefit card
- Prior authorization documentation (optional)
- Release of medical records forms (optional)
4.2 System Requirements
- Electronic Health Record (EHR) access
- Insurance eligibility verification portal
- PBM interface connectivity
- Risk assessment calculation tool
5. Main Procedure
5.1 Insurance Validation Protocol
5.1.1 Primary Insurance Verification
- Access payer portal using provider credentials
- Verify coverage status and effective dates
- Document benefit levels and copayment requirements
- Confirm network participation status
- Record authorization requirements
5.1.2 Prescription Benefit Validation
- Query PBM database for current coverage
- Verify formulary compliance parameters
- Document prior authorization requirements
- Confirm specialty pharmacy protocols
- Record copayment structure
5.2 Risk Assessment Protocol
5.2.1 Lifestyle Risk Evaluation
Patient Intake and Registration SOP - Part 2
- Calculate smoking risk index (Never=0, Former=1, Current=2)
- Assess alcohol consumption (None=0, Occasional=1, Moderate=2, Heavy=3)
- Evaluate exercise patterns (5+ times=-1, 3-4 times=0, 1-2 times=1, None=2)
- Compute aggregate lifestyle score
5.2.2 Clinical Risk Assessment
- Review surgical history chronology
- Evaluate chronic condition severity
- Calculate comorbidity interaction score
- Generate weighted risk factor index
5.3 Patient Registration Protocol
5.3.1 Eligibility Confirmation
- Verify insurance_validation status = "valid"
- Confirm prescription_insurance_validation = "valid"
- Validate life_style_risk_level within parameters
- Assess overall_risk_level compatibility
5.3.2 Pharmacy Network Verification
- Query preferred pharmacy database
- Verify pharmacy_check status = "yes"
- Confirm pharmacy network participation
- Document special handling requirements
6. Output
6.1 Required Documentation
The procedure generates a verification report (json) with all results.
These include insurance validation check, prescription validation check,
pharmacy check, life style risk, overall risk, and registration status.
All outputs must be archived in compliance with regulatory requirements.
6.2 System Updates
- Active patient status in EHR
- Documented risk levels
- Verified insurance records
- Confirmed pharmacy selection
- Completed registration status
F.2 Sample Patient Data (Synthetically Generated)
Table 5: Sample Patient Registration Data (First Five Records)
Patient ID First Name Last Name DOB Age Insurance Smoking Alcohol Exercise Insurance Valid Risk Level Reg
P100001 John1 Dove 1980-05-15 43 Blue Cross Never Occasional 3-4 times valid low success
P100002 John2 Dove 1995-08-22 28 Aetna Current Moderate 1-2 times valid medium success
P100003 John3 Dove 1972-03-10 51 United Former Heavy None valid high failure
P100004 John4 Dove 1990-11-30 33 Cigna Never None 5+ times valid low success
P100005 John5 Dove 1965-07-25 58 Medicare Former Occasional 1-2 times invalid high failure
F.3 Tool Specifications
Listing 9: Tool Specifications in JSON Format
1
2[
3 {
4 "toolSpec": {
5 "name": "calculateLifestyleRisk",
6 "description": "Calculates a patients lifestyle risk level based on smoking status, alcohol consumption, and exercise frequency, following the standardized risk assessment protocol defined in the SOP.",
7 "inputSchema": {
8 "json": {
9 "type": "object",
10 "properties": {
11 "patient_id": {
12 "type": "string",
13 "description": "Unique identifier for the patient",
14 "pattern": "^P[0-9]{9}\\$",
15 "example": "P123456789"
16 },
17 "smoking_status": {
18 "type": "string",
19 "enum": [
20 "Never",
21 "Former",
22 "Current"
23 ],
24 "description": "Patients current smoking status, used to calculate smoking risk index"
25 },
26 "alcohol_consumption": {
27 "type": "string",
28 "enum": [
29 "None",
30 "Occasional",
31 "Moderate",
32 "Heavy"
33 ],
34 "description": "Patients level of alcohol consumption frequency"
35 },
36 "exercise_frequency": {
37 "type": "string",
38 "enum": [
39 "None",
40 "1-2 times",
41 "3-4 times",
42 "5+ times"
43 ],
44 "description": "Weekly frequency of patients exercise activities"
45 }
46 },
47 "required": [
48 "patient_id",
49 "smoking_status",
50 "alcohol_consumption",
51 "exercise_frequency"
52 ],
53 "additionalProperties": false
54 }
55 }
56 }
57 },
58 {
59 "toolSpec": {
60 "name": "verifyPharmacy",
61 "description": "Validates pharmacy network participation and verifies pharmacy details as part of the patient registration process",
62 "inputSchema": {
63 "json": {
64 "type": "object",
65 "properties": {
66 "patient_id": {
67 "type": "string",
68 "description": "Unique identifier for the patient",
69 "pattern": "^P[0-9]{9}\\$",
70 "example": "P123456789"
71 },
72 "preferred_pharmacy_name": {
73 "type": "string",
74 "description": "Full legal name of the preferred pharmacy",
75 "minLength": 1,
76 "maxLength": 100,
77 "example": "CVS Pharmacy"
78 },
79 "preferred_pharmacy_address": {
80 "type": "string",
81 "description": "Complete street address of the pharmacy including city, state, and zip code",
82 "minLength": 5,
83 "maxLength": 200,
84 "example": "123 Main St, Boston, MA 02108"
85 },
86 "pharmacy_phone": {
87 "type": "string",
88 "description": "Contact phone number for the pharmacy in XXX-XXX-XXXX format",
89 "pattern": "^[0-9]{3}-[0-9]{3}-[0-9]{4}\\$",
90 "example": "555-123-4567"
91 }
92 },
93 "required": [
94 "preferred_pharmacy_name",
95 "preferred_pharmacy_address",
96 "pharmacy_phone"
97 ],
98 "additionalProperties": false
99 }
100 }
101 }
102 }
103]